If the equilibrium constant for a two electron redox reaction at 198 oC is 1500, calculate the  ΔG° and E0red.
Answer #1

∆G° = -{8.314*471*ln(1500)} = -28.638 kJ/mol

red = -∆G°/nF = 28638/(2*96485) = 0.1484 V